Chuck end Jackie stand on separate carts, both of
Answer #1

(a) Here we will have to calculate how fast Chuck and the ball are moving away from each other. So, we will have to calculate the relative speed between the two

u = vc + vb

(b) initial momentum of Chuck, cart and the ball pi = 0, as they all are at rest

Final momentum of Chuck, cart and the ball pf = 0 + mcart (-vc) + mball vb

pf = -mcart vc + mball vb

from part (a) we have vc = u - vb

pf = -mcart (u - vb) + mball vb

pf = -mcart u + (mcart + mball ) vb

Now, by conservtion of momentum pi = pf

0 = -mcart u + (mcart + mball ) vb

vb = (mcart u) / (mcart + mball )

(c) From part (a) we have vc = u - vb

Using value of vb from part (b), we get

vc = u - { (mcart u) / (mcart + mball ) }

vc = (mball u) / (mcart + mball )